1. Suppose that a seller decreases price for his good by 5 percent show how total revenue changes when elasticity of demand =0.5
2. Assume that good A and good B are related goods and QB=1691-400PB+6PA-6Y. Suppose that PB=0.1 dollar, PA=0.3 Dollar and Y(income)=10 dollar then compute.
A. Price elasticity of demand for good B
B. Income elasticity of demand for good
B and explain nature of the good
C. Cross price elasticity of demand for
good B and explain nature of the good
whether they are substitute,
complementary, or unrelated.
1
Expert's answer
2019-12-16T09:54:00-0500

1. If a seller decreases price for his good by 5 percent, then his quantity demanded will decrease by 2.5 percent, so the total revenue will increase by 1.05/1.025 - 1 = 0.0244 or 2.44 percent when elasticity of demand =0.5.

2. QB = 1691 - 400×0.1 + 6×0.3 - 6×10 = 1646.8 units.

A. Price elasticity of demand for good B is: Ed = -400×0.1/1646.8 = -0.024, so the demand is inelastic.

B. Income elasticity of demand for good B is:

Ei = -6×10/1646.2 = -0.036, so the good is inferior.

C. Cross price elasticity of demand for

good B is:

Ecp = 6×0.3/1646.2 = 0.001, so the goods are substitutes.